OpenStudy (anonymous):

find f(1/14)+f(2/14)+...+f(13/14)

OpenStudy (anonymous):

\[\large f(x)=\frac{4^x}{4^x+2}\]

OpenStudy (anonymous):

I think we have to solve f(x) more..

OpenStudy (anonymous):

1/14+13/14=1

OpenStudy (anonymous):

Yes... Like wise 2/14 + 12/14 = 1

OpenStudy (anonymous):

simplify f(x)+f(1-x)

OpenStudy (asnaseer):

yes - I believe @JohnHanShanghai has the correct approach.

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

see i want to start with an example